OpenStudy (anonymous):

The base of a triangle is 3 cm less than 4 times the height of the triangle. The area of the triangle is 26 cm2. Find the base and height of the triangle.

OpenStudy (anonymous):

the base becomes 4h-3 now use the formulae for area=1/2base*height and solve for h

OpenStudy (anonymous):

let the height of the triangle be x cm then its base is 4x-3 now area of the triangle is given by 1/2 * base * height hence 1/2 *(4x-3)*x= 26

OpenStudy (anonymous):

whats the answer for them tho?

OpenStudy (anonymous):

it is 4 u will get two values of h one positive and one negetive consider the positive value as height is never negetive and b=4h-3=13

OpenStudy (anonymous):

4x^2 -3x -52 =0 4x^2-16x+13x-52=0 4x(x-4)+13(x-4)=0 (x-4)(4x+13)=0 thus x-4 =0 give x=4 and 4x+13 give x=-14/3 (not possible ) hence x=4 thus height of the triangle is 4 cm and the base is 4*4 -3=13cm

OpenStudy (anonymous):

so h=4 and b=13

OpenStudy (anonymous):

ok thanks alot :)

OpenStudy (anonymous):

ur welcome